OpenStudy (anonymous):

subtract. Simplify by collecting like radicals terms if possible, assuming that all expressions under radicals represent non-neg numbers (see comments for work)

OpenStudy (anonymous):

\[\sqrt[3]{1250x}-\sqrt[3]{10x ^{4}}\]

OpenStudy (anonymous):

\[5\sqrt[3]{10x}-\sqrt[3]{10x ^{4}}\]

OpenStudy (anonymous):

factor out x^3 and cube root 10

OpenStudy (anonymous):

\[\sqrt[3]{10}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you write the whole equation?

OpenStudy (anonymous):

\[\sqrt[3]{10}=10000\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no

OpenStudy (anonymous):